# Aerial Vision-Language Navigation with a Unified Framework for Spatial, Temporal and Embodied Reasoning

### v1

@@ IV-C1 Effect of Auxiliary Tasks
IV-C2 Effect of History Representation Strategies
TABLE IV: Comparison of different history representation strategies on AerialVLN-S Val-Unseen split.
@@ IV-C2 Effect of History Representation Strategies
Figure 9: Success rate across different trajectory lengths on the Val-Unseen split. Long-Horizon Uniform Sampling maintains stable performance across episodes of varying lengths.
We compare three history representation strategies. (1) Current-Only, where the model relies solely on the latest observation, often leading to unstable behaviors such as repeated fine-grained heading corrections. (2) Short-Term Memory Bank, which maintains a fixed-size memory bank following a first-in-first-out (FIFO) policy, keeping only the most recent observations and discarding earlier ones [gao2025openfly]. This provides short-range temporal cues while discards earlier information essential for global navigation. (3) Long-Horizon Uniform Sampling, which uniformly samples a small set of frames from the entire past trajectory to provide long-range visual context. Table IV compares these history representation strategies. Among them, Long-Horizon Uniform Sampling ($k=8$) achieves the best overall performance. By drawing observations uniformly from entire past trajectory, this strategy preserves both short-term motion cues and long-range scene information. As a result, it provides a more reliable temporal context than relying solely on the current frame or restricting history to a short FIFO queue. We compares these strategies using the NVILA-2B model for computational efficiency. The larger 8B variant follows the same trend and achieves consistently higher accuracy.
@@
IV-C3 Effect of Action Merging, Keyframe Selection amd Lable Reweighting
Table V reports the effect of action merging (AM), keyframe selection (KS), and label reweighting (LR). For computational efficiency, this ablation uses the NVILA-2B model without auxiliary tasks, and all variants adopt a long-horizon uniform sampling of 8 history frames. Enabling action merging leads to immediate gains, as consolidating micro-actions into coherent motion segments produces cleaner supervisory signals compared to raw fragmented steps. Excessively large merge lengths slightly degrade performance, likely because they skip important visual cues and reduce the model’s ability to adapt to local changes. We therefore set the maximum merge length to 3. Adding keyframe selection further improves navigation accuracy. Selecting the boundaries of merged actions as keyframes provides a more structured trajectory history in which visual transitions align with meaningful motion changes. This reduces temporal redundancy while preserving observations most relevant to spatial landmarks. Finally, incorporating label reweighting achieves the best overall performance. By balancing the contribution of long merged segments and short corrective actions, it prevents the model from overfitting to frequent but low-impact motions and strengthens action-level supervision. Overall, the combined use of AM, KS, and LR yields substantial improvements in SR and SDTW, demonstrating that structuring the action sequence and rebalancing its supervision are both crucial for reliable aerial navigation.
TABLE V: Ablation study on the effect of action merging (AM) , keyframe selection (KS) and label reweighting (LR) on AerialVLN-S Val-Seen split. All of them consistently improve performance across all metrics.
@@ IV-C3 Effect of Action Merging, Keyframe Selection amd Lable Reweighting
IV-D Failure Case Analysis
Although our method achieves strong performance and outperforms competitive baselines, it is still susceptible to several typical failure modes. To better understand the remaining challenges and guide future improvements, we further analyze its most common failure cases below. A common failure mode arises from ambiguous instructions that provide only a sequence of low-level actions without any accompanying landmark references (e.g., “rotate, go straight, turn left, turn right, hold on”). Such instructions lack grounding anchors that would normally guide the agent in determining where a directional change or stop should occur. Under this setting, the agent receives no observable cues to align the action sequence with the environment, making it difficult to infer the intended spatial context. Another failure mode arises from the intrinsic limitations of single-view RGB perception in complex outdoor environments. When the agent lacks sufficient observational coverage, such as when critical cues briefly appear at the edge of view, the model may fail to terminate at the correct moment or overshoot the intended point, resulting in unnecessary backtracking or redundant movement. This behavior reflects the inherent difficulty of aligning long-horizon instructions with first-person observations. Enhancing viewpoint robustness, increasing the diversity of training trajectories, or exposing the model to more error-recovery scenarios in simulation could help mitigate such issues.
V Conclusion
@@
[Uncaptioned image] Huilin Xu (Student Member, IEEE) received the B.E. degree with honors from the School of Information Science and Engineering, Fudan University, Shanghai, China, in 2021. She is currently pursuing a Ph.D. degree at the Key Laboratory of Information Science of Electromagnetic Waves (MoE), Fudan University. Her research interests focus on visual reasoning, vision-based reinforcement learning, and robotic learning.
[Uncaptioned image] Zhuoyang Liu (Student Member, IEEE) received the B.E. degree from Wuhan University, Wuhan, Hubei, China, in 2020. She is currently pursuing the Ph.D. degree in electromagnetic science with the Key Laboratory of Information Science of Electromagnetic Waves, Fudan University, Shanghai, China. Her research interests include signal processing, Dual-Function Radar-Communication systems, reconfigurable intelligence surfaces, and the combination of artificial intelligence and electromagnetic waves.
[Uncaptioned image] Yixiang Luomei (S’18-M’22) received the B.E. and M.E. in Electronic and Communication Engineering from Xidian University, Xi’an, China, and the Ph.D in Electronic Information Engineering from Fudan University, Shanghai, China, in 2014, 2018 and 2022, respectively. From 2022 to 2023, she was a Research Fellow with the National University of Singapore, Singapore. Since 2013, she has been a Research Assistant of the MoE Key Lab for Information Science of Electromagnetic Waves, School of Information Science and Technology, Fudan University, Shanghai, China. She has published more than 10 papers in peer-reviewed journals, among some conference papers and patents. She is currently working on intelligent processing of electromagnetic signals, UAV SAR imaging, and ISAC sensing.
[Uncaptioned image] Feng Xu (Member, IEEE) received the B.E. degree (Hons.) in information engineering from Southeast University, Nanjing, China, in 2003, and the Ph.D. degree (Hons.) in electronic engineering from Fudan University, Shanghai, China, in 2008. From 2008 to 2010, he was a Post-Doctoral Fellow with the National Oceanic and Atmospheric Administration (NOAA) Center for Satellite Applications and Research, Camp Springs, MD, USA. From 2010 to 2013, he worked with Intelligent Automation Inc., Rockville, MD, USA, and NASA Goddard Space Flight Center, Greenbelt, MD, USA, as a Research Scientist. In 2012, he was selected for China’s Global Experts Recruitment Program and subsequently returned to Fudan University, in 2013, where he is currently a Professor and the Vice Dean of the School of Information Science and Technology and the Vice Director of the electromagnetic waves (MoE) Key Laboratory for Information Science of Electromagnetic Waves. His research interests include electromagnetic scattering modeling, SAR information retrieval, and radar system development. Dr. Xu is a Topic Associate Editor of the IEEE TRANSACTIONS OF GEOSCIENCE AND REMOTE SENSING and was an Associate Editor of the IEEE GEOSCIENCE AND REMOTE SENSING LETTERS (2014–2021). He is the Founding Chair of the IEEE GRSS Shanghai Chapter and an IEEE GRSS AdCom Member.

### v2

@@ IV-D Ablation Study
IV-D1 Effect of Auxiliary Tasks
We conduct ablation experiments to evaluate the effect of the two auxiliary tasks: Spatial Perception (SP) and Trajectory Reasoning (TR) in Table VI. Without auxiliary supervision, the policy shows limited spatial grounding and produces the lowest SR and SDTW on both splits. SP leads to consistent gains, confirming that geometric cues improve local action grounding. Incorporating TR further boosts performance, with the largest improvements on unseen environments, where long-horizon cues are critical for handling novel layouts. These results demonstrate that the two objectives are complementary: SP enhances local spatial understanding, while TR strengthens temporal reasoning. Their combination yields a more reliable navigation policy and supports the design of our multi-task framework, reinforcing our goal of unifying spatial, temporal, and embodied reasoning within a single model.
IV-D2 Effect of History Representation Strategies
TABLE VII: Comparison of different history representation strategies on AerialVLN-S Val-Unseen split.
@@ IV-D2 Effect of History Representation Strategies
Figure 11: Success rate across different trajectory lengths on the AerialVLN-S Val-Unseen split. Long-Horizon Uniform Sampling maintains stable performance across episodes of varying lengths.
We compare three history representation strategies. (1) Current-Only, where the model relies solely on the latest observation, often leading to unstable behaviors such as repeated fine-grained heading corrections. (2) Short-Term Memory Bank, which maintains a fixed-size memory bank following a first-in-first-out (FIFO) policy, keeping only the most recent observations and discarding earlier ones [12]. This provides short-range temporal cues while discards earlier information essential for global navigation. (3) Long-Horizon Uniform Sampling, which uniformly samples a small set of frames from the entire past trajectory to provide long-range visual context. Table VII compares these history representation strategies. Among them, Long-Horizon Uniform Sampling ($k=8$) achieves the best overall performance. By drawing observations uniformly from entire past trajectory, this strategy preserves both short-term motion cues and long-range scene information. As a result, it provides a more reliable temporal context than relying solely on the current frame or restricting history to a short FIFO queue. We compares these strategies using the NVILA-2B model for computational efficiency. The larger 8B variant follows the same trend and achieves consistently higher accuracy.
@@
TABLE VIII: Ablation study on the effect of action merging (AM) , keyframe selection (KS) and label reweighting (LR) on AerialVLN-S Val-Seen split. All of them consistently improve performance across all metrics.
@@
IV-D3 Effect of Action Merging, Keyframe Selection, and Label Reweighting
Table VIII reports the effect of action merging (AM), keyframe selection (KS), and label reweighting (LR). For computational efficiency, this ablation uses the NVILA-2B model without auxiliary tasks, and all variants adopt a long-horizon uniform sampling of 8 history frames. Enabling action merging leads to immediate gains, as consolidating micro-actions into coherent motion segments produces cleaner supervisory signals compared to raw steps. Adding keyframe selection further improves navigation accuracy. Selecting the boundaries of merged actions as keyframes provides a more structured trajectory history in which visual transitions align with meaningful motion changes. This reduces temporal redundancy while preserving observations most relevant to spatial landmarks. Finally, incorporating label reweighting achieves the best overall performance. By balancing the contribution of long merged segments and short corrective actions, it prevents the model from overfitting to frequent yet trivial motions, thereby strengthening action-level supervision.
We further ablate the maximum number of consecutive actions merged (i.e., the merge step) and report the results in Table IX. As shown, a merge step of 3 achieves the best performance, while a merge step of 1 (i.e., no merging) retains the original fragmented action sequence and leads to lower navigation performance. Excessively large merge steps (e.g., 6) slightly degrade performance, likely because they skip important visual cues and reduce the model’s ability to adapt to local changes. We therefore set the maximum merge length to 3. Keyframes are then selected at the boundaries of these refined segments, ensuring sufficient temporal coverage and preserving landmarks that may appear along extended straight paths. Overall, the combined use of AM, KS, and LR yields substantial improvements in SR and SDTW, demonstrating that structuring the action sequence and rebalancing its supervision are both crucial for reliable aerial navigation.
TABLE IX: Ablation on the maximum number of consecutive actions merged evaluated on AerialVLN‑S val‑seen split. The best performance is achieved with a merge step of 3.
[…]
1 90.2 1.8 20.5 0.7
3 87.1 5.7 28.9 1.9
6 89.5 5.1 25.3 1.4
TABLE X: Quantitative failure case analysis on the AerialVLN-S val-unseen split. Note that $\dagger$ indicates our method without auxiliary tasks. $\Delta$ denotes the difference in episode counts between the proposed model and the variant trained without auxiliary tasks.
Failure Type Ours Ours $\Delta$
Perception-related 194 204 10-10
Long-horizon drift 124 135 11-11
Collision 142 150 8-8
Stop failure 107 90 +17+17
IV-E Failure Case Analysis
Although our method achieves strong performance and outperforms competitive baselines, it is still susceptible to several typical failure modes. To better understand the remaining challenges and guide future improvements, we further analyze its most common failure cases below. A common failure mode arises from ambiguous instructions that provide only a sequence of low-level actions without any accompanying landmark references (e.g., “rotate, go straight, turn left, turn right, hold on”). The absence of such grounding references makes it difficult for the agent to determine where in the environment each action should be executed, leading to misalignment between the instruction and the visual context. We manually identify 21 severely underspecified instructions whose action sequences lack explicit landmark references. On this subset, both the proposed model and the variant without auxiliary tasks fail. This observation is consistent with previous work[45] as the absence of explicit landmark references prevents the agent from grounding the action sequence to specific locations in the environment, leading to inherently unreliable decision-making. Excluding these cases, we analyze the remaining episodes using trajectory-level metrics and conduct a quantitative analysis on the AerialVLN-S val-unseen split. Failure cases are categorized into stop failures (OSR=1 with incorrect termination), collision failures (OSR=0 with obstacle contact), long-horizon drift (OSR=0 with low trajectory similarity measured by nDTW), and perception-related failures (remaining cases exhibiting partial trajectory alignment but incorrect intermediate grounding). As summarized in Table X, the auxiliary tasks reduce both long-horizon drift and perception-related failures compared to the model trained without auxiliary supervision. This suggests improved global alignment over extended trajectories and more stable visual grounding under monocular observation. Collision failures are slightly reduced. In contrast, stop failures increase, indicating that the auxiliary tasks enable the agent to reach the goal region more frequently while termination decisions remain imperfect. Overall, the results show that the auxiliary tasks mitigate specific failure modes, particularly long-horizon drift and perception-related errors, by strengthening alignment between visual perception, textual language, and embodied action.
V Conclusion and Discussion
@@
Despite the encouraging results, several limitations remain. First, the current framework relies on passive monocular observations, which may lead to incomplete environmental coverage under restricted field-of-view settings. Second, performance can degrade under ambiguous instructions that lack explicit landmark references. Third, long-horizon trajectory alignment remains challenging when small prediction errors accumulate over time. Future work will explore active viewpoint control to expand perceptual coverage, ambiguity-aware instruction clarification to improve grounding under underspecified commands, and hierarchical or staged planning strategies to enhance long-range consistency. Moreover, aerial VLN method could benefit from lightweight obstacle-aware planning modules that operate alongside the navigation policy.